OpenStudy (anonymous):

how do I solve the following system of equations? x=15+6y 9x+3y=21

OpenStudy (anonymous):

plug the first equation into the second one's x

OpenStudy (anonymous):

Ive gotten this far 9x+3y=21 9(15+6y)+3y=21

OpenStudy (anonymous):

9(15+6y) +3y=21 and solve

OpenStudy (anonymous):

okay, now distribute

OpenStudy (anonymous):

Ik you distribute the 9 which would be 135+57y=21

OpenStudy (anonymous):

good then subtract 135 from 21

OpenStudy (anonymous):

so you get 57y= -114. divide the -114 by 57 and voila! you get your answer

OpenStudy (anonymous):

k I did it right then

OpenStudy (anonymous):

hope that helped

OpenStudy (anonymous):

but what will I put in for x when I check my answer

OpenStudy (anonymous):

so y=-2 right? take the -2 and plug it into y in either equation to get x

OpenStudy (anonymous):

with it being 9x+3y=21

OpenStudy (anonymous):

in both equations, x=3

OpenStudy (anonymous):

how did you get that ?

OpenStudy (anonymous):

sorry slowly getting the hang of this

OpenStudy (anonymous):

its ok:)

OpenStudy (anonymous):

because u used bothe equations to find y, they will also share the same x. so when y is plugged into one equation, x will equal 3. double check with the other equation and we also get 3 for x

OpenStudy (anonymous):

is it clear now? or do you have more questions?

OpenStudy (anonymous):

still trying to figure out how you get 3?

OpenStudy (anonymous):

ok, we take the -2 that is y. you pick either equation, it does not matter. x=15+6(-2). x=15+(-12). x=3

OpenStudy (anonymous):

oh ok I get it I am very sorry

OpenStudy (anonymous):

it ok, happy to help :)

OpenStudy (anonymous):

thank you for be patient with me

OpenStudy (anonymous):

no prob.